OpenStudy (anonymous):

Which equation describes the translation of the equation? Y=IxI 5 units left

OpenStudy (tyteen4a03):

This graph should help: 

OpenStudy (anonymous):

A: y=IxI+ 5 B: Y=Ix+5I C: Y=I-5XI D: Y-5=IxI